OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b? A. 4,1 B. -19,40 B. 20,1 C. 30,9 D. -8,-29

OpenStudy (anonymous):

This is what I got ifference of two numbers is 21: a - b = 21 difference of 5 times a and 2 times b is 18: 5a - 2b = 18 we have a system of equations: a = 21 + b 5a - 2b = 18 5(21 + b) - 2b = 18 105 + 5b - 2b = 18 3b = -87 b = -29 Now that we have b, solve for a: a = 21 + b a = 21 + (-29) a = -8 Your numbers are: -8 and -29

OpenStudy (anonymous):

Thank you so much ! (:

OpenStudy (anonymous):

You are very welcome! :)

OpenStudy (anonymous):

Do you think you can help me with another question?